OpenStudy (anonymous):

Solved the system by substitution.if the system is inconsistent and has no solution,state this .if the system is dependent,write the form of the solution for any real number x. x+y=9000 0.06x+0.15y=1215

hero (hero):

Hint: y = 9000 - x

OpenStudy (anonymous):

Can you explain steps and steps ,so as I can be familiar with it

OpenStudy (anonymous):

Substitute the equation Hero gave you into the other equation 0.06x+0.15(9000-x)=1215 solve for x

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

x=1214

OpenStudy (anonymous):

I divide 1215-0.15

OpenStudy (anonymous):

I didnt get it right

OpenStudy (anonymous):

I dont think so 0.06x+1350-0.15x=1215 YOU need to subtract from 1215

OpenStudy (anonymous):

1215- (0.06x+1350)=135.06

OpenStudy (anonymous):

.06x-0.15x=1215-1350

OpenStudy (anonymous):

-0.09x=-135

OpenStudy (anonymous):

Are you still here

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so you will subtract -0.09-135

OpenStudy (anonymous):

-0.09x=-135

OpenStudy (anonymous):

I dont think so 0.06x+1350-0.15x=1215 YOU need to subtract from 1215

OpenStudy (anonymous):

-0.09x=-135 divide both sides by -0.09

OpenStudy (anonymous):

I am confused not good in math

OpenStudy (anonymous):

-0.09x/-0.09=-135/-0.09

OpenStudy (anonymous):

x=1500 After you have x, substitute back into the equation to get y. y=9000-1500 y=7500 You can check to see if values work into the other equation. If it doesn't, you've messed up. check: 0.06(1500)+0.15(7500)=1215 90+1125=1215 1215=1215 our x and y values are right

OpenStudy (anonymous):

Does that help... the only way to get better is to keep practicing so you're doing the right thing...asking questions and hanging in there

OpenStudy (anonymous):

I will try ,I think I need to start from simplest to complex

OpenStudy (anonymous):

is that the final answer to the equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(1500,7500)

OpenStudy (anonymous):

So is the equation system a inconsistent and has no solution or depedent or independent

OpenStudy (anonymous):

dependent

OpenStudy (anonymous):

what about the first one that you solved

OpenStudy (anonymous):

which one was that...I surf the site

OpenStudy (anonymous):

this is the one am talking about 9x-6y=19 -x+8y=4 x=8y-4 9(8y-4)-6y=19 72y-36-6y=19 66y=55 y=5/6

OpenStudy (anonymous):

So is the equation system a inconsistent and has no solution or depedent or independent

OpenStudy (anonymous):

independent there is only one point on a graph where these two lines will intersect